Give two equations using x and y that have (1,2) as a solution

Found 2 solutions by ewatrrr, ikleyn:
Answer by ewatrrr(24785)


Hi
Give two equations using x and y that have (1,2) as a solution 
 1)  y-2 = 2(x-1)   0r  y = 2x 

 2) y-2 = (-1/2)(x-1)  0r  y = -.5x + 5/2



Wish You the Best in your Studies.


Answer by ikleyn(53751)

.

One equation x = 1.

Another equation y = 2.